OpenStudy (anonymous):

i have no idea how to set this up Yuka buys 6 pounds of apples at $0.50 a pound. She also buys 6 pounds of bananas. Altogether, the fruit costs $12.00. What is the cost per pound for the bananas? A. $2.00 B. $1.00 C. $1.50 D. $0.50

OpenStudy (anonymous):

ok so she buys 12 pounds in total, and half of that are apples. you know the price of the apples. if its .50 per pound, whats 6x.50? then take that number, and subtract it from 12. then take that result and divide it by 6 and you will get your answer

OpenStudy (anonymous):

so it would be 1.50?

OpenStudy (anonymous):

yep! good job :)

OpenStudy (anonymous):

nah, you just have to read it one part at a time

OpenStudy (anonymous):

i do is just that all the words confuse me kinda thats why i i dont like word problems thats also the reason why im taking 9th grade math when im in 10th grade

OpenStudy (anonymous):

oh I see. well I can simplify this down for you if you want so you can use it for future word problems? word problems are my forte :P

OpenStudy (anonymous):

all i need help with simplifying world problems and thanks that would be very nice of you :)

OpenStudy (anonymous):

ok so Yuka buys 6 pounds of apples at $0.50 a pound. She also buys 6 pounds of bananas. Altogether, the fruit costs $12.00. What is the cost per pound for the bananas? If equations are easier for you, use the information to make an equation. theres 12 pounds total (6+6) so that will be your total (x+x=12) and you know 1 pound of apples is .50 you also know she bought 6 pounds of bananas. so the equation qould be 1.50x6+bx6=12 (b=bananans, and there were 6 pounds total) so solving that equation should get you 1.50 (Ithink you need to divide 9/6 not 6/9) I hope that helps :/ I'm not that good at explaining

OpenStudy (anonymous):

i dont say your not good because you actually helped me :)

OpenStudy (anonymous):

aw ok c: let me know if you need any more word problem help

OpenStudy (anonymous):

i have another one At her job, Mary works 4 hours on Monday, 3 hours on Wednesday, and 6 hours on Saturday. Her boss adds a $10 bonus to her paycheck, which then totals $120.50. What is Mary's hourly pay? A. $8.50 B. $9.27 C. $5.24 D. $8.05 wouldnt you do 120.50-10=110.5 and then add the hours together and divide 110.5 by the hours?

OpenStudy (anonymous):

hmm I would subtract 10 first because thats her bonus. then I'd add the hours together.. so 13 hours and theres 110.50 total, then divide.. the answer is A, 8.50 sooo YES you are correct :D *throws confetti*

OpenStudy (anonymous):

omg!!!! im like so proud lol

OpenStudy (anonymous):

thank u!!!!!

OpenStudy (anonymous):

:D your welcome <3

OpenStudy (anonymous):

A vine 6 cm tall is planted in a garden. It grows 4 centimeters a day. How many days will it take to reach 122 cm in height? A. 32 B. 30.5 C. 29 D. 27 for this one do you do 122 divided by 4 and you get your answer which is 30.5? or do you have to do something with the 6cm?

OpenStudy (anonymous):

ok so its already 6 cm, so you would subtract that from the total (always do this) so thats 116 then take 116/4 (because that gives you the days) so yea you have to do something with the 6, because if its already 6 you dont need it in the equation. (theres really 3 kinds of info in a problem, the starter, the total, and the variable, the total needs to have the started subtracted form it, and the variable is what you use to figure out the answer)

OpenStudy (anonymous):

i thought you didnt need to use the 6 because it was already that tall, but when you solve that the answer is 29 right?

OpenStudy (anonymous):

yep c: tbut yea if somehting is already that tall, or like with the bonus theres an extra 10 bucks, and with the apples it was the .50, but theres always an extra something that you need to take away from the total.

OpenStudy (anonymous):

ohhhh okayyy i get that partt now ,umm now i really need your help understanding this because im doing Translate and Solve Written Statements

OpenStudy (anonymous):

I can help you with them :) just write the problem and write what you did to solve and if it's wrong I'll help you with what you got wrong ok c:

OpenStudy (anonymous):

okayy :D

OpenStudy (anonymous):

Twice the difference between a number and 4 is twenty. Find the number. A. 8 B. 9 C. 12 D. 14 so when it says twice the difference does it mean to do 2x+4=20?

OpenStudy (anonymous):

well dfference is subtraction (the difference in 5 and 4 is one) I honestly don't know this one but since I'm a cheater I googled it and yes its 2x+4=20 xD

OpenStudy (anonymous):

lol nice one so the answer would be 8?

OpenStudy (anonymous):

yep :D

OpenStudy (anonymous):

:D In a change jar, Tom has nickels and dimes. The number of nickels is two times the number of dimes and the total amount of money in the jar is $1.20. How many dimes are in the jar? A. 6 B. 12 C. 10 D. 24 so this one is kinda hard would it be 2d=1.20?

OpenStudy (anonymous):

and it would be 6?

OpenStudy (anonymous):

ok so theres 2times more nickels then dimes...and 1.20 total I believe it is 6. so if there were 6 dimes that would be 60. that would mean there would HAVE to be 12 nickels and that equals to 60 as well so yes :D you're right

OpenStudy (anonymous):

woa actually ive never done this good i guess it helps having a person to tell what you think and they can say if ur right or not The length of a rectangle is five centimeters more than its width, w. If the perimeter of the rectangle is 26 centimeters, what is the area of the rectangle, in square centimeters? A. 104 B. 676 C. 36 D. 26 so for this one would it be 5w=26 (idk this one confuses me)

OpenStudy (anonymous):

so perimeter is the sum of all 4 sides. and the l and 5 more than w so if the sum is 26 aw man this is hard well the sides are 4 and 9, so 4*9 is 36 I dont know how I came up with the answer but the answer is 36 @_@ I just plugged in numbers until I got 4 and 9 and that gives you the answer

OpenStudy (anonymous):

lol that was the same thing i said this is hard but is basically plugging in random numbers?

OpenStudy (anonymous):

well it had to equal 26, and you should start with a low number so I tried 3 and 8 (remember one is 5 mroe than the toehr) and that was really close so I bumped it up one and got it :P sometimes you have to do that if it's hard

OpenStudy (anonymous):

yeah i do that sometimes but i didnt know you could do that for this problem

OpenStudy (anonymous):

you can do it for any problem that involves a variable really

OpenStudy (anonymous):

Eight less than a number doubled is twenty more than the number. What is the number? A. 28 B. 6 C. 14 D. 36 for this one is 8-n= ugh im stuck

OpenStudy (anonymous):

well i got to go

OpenStudy (anonymous):

this one just plug in each answer until it fits so 6... double that and its 12 and eight less is not 20 more then the number

OpenStudy (anonymous):

hey can u help me today?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

I have to go soon but I can try

OpenStudy (anonymous):

alright i only have like 2 questions

OpenStudy (anonymous):

The area of a triangle is 22 , with a base of 8 yards. What is the height, in yards? A. 5.5 B. 11 C. 2.75 D. 20 for this one i set up the equation 1/2b(base) times (height) and i plugged in 8 into the base and then got 4 but i tried dividing that buy 22 but it didnt give me a right answer

OpenStudy (anonymous):

it's 5.5 yards Area = 1/2 * Base * Height 22 = 1/2 * 8 * Height Height = 22/4 = 5.5 yd

OpenStudy (anonymous):

with geometry equations you just have to remember a ton of formulas and I think thats what you got wrong really, is that it wasnt the right formula (:

OpenStudy (anonymous):

im not good remembering stuff honestly